Find the accumulated amount A if the principal P = $11,000 is invested at the interest rate of r = 5% per year for t = 5.5 years, compounded quarterly.
Answer by rfer(16322)

A=11000(1+0.05/4)^4*5.5
A=11000(1.0125)^22
A=11000*1.31429=$14,457.00